OpenStudy (anonymous):

The ratio of fruit juice to ginger ale in a punch is 2 to 3. How much ginger ale is needed to make 25 liters of punch?

OpenStudy (anonymous):

16 and 2/3 liters

OpenStudy (anonymous):

2 / 3 = x / 25 3x = 50 x = 50/3 = 16 2/3